# Efficient Semi-Supervised Adversarial Training via Latent Clustering-Based Data Reduction

### v2

@@ Appendix C Additional Experiments
C-A Hyperparameter Sensitivity Analysis Details
In Section V-C, we present a detailed analysis of how accuracies are influenced by different hyperparameter values. In these experiments, we varied one hyperparameter at a time while keeping the others constant. Figure 4a illustrates how robust accuracy on the CIFAR-10 dataset changes with $\beta$ values when $20\%$ additional data is incorporated from the 11M data samples generated by the DDPM model, using a per-batch ratio of $0.3$. Figure 4b shows the variation in robust accuracy for the SVHN dataset as $\beta$ changes, with $10\%$ drawn from the SVHN extra dataset and a per-batch ratio of $0.5$.
In Figure 7a, we present how SSAT’s robust accuracy is affected on CIFAR-10 by varying $\gamma$ when $20\%$ extra data from the 11M generated DDPM samples is used, with $\beta$ fixed at $0.4$. Similarly, Figure 7b displays the effect of varying the per-batch ratio on robust accuracy for the SVHN dataset, using $10\%$ data from the SVHN extra dataset and fixing $\beta$ at $0.6$. In addition, Figures 4c and 4d explore how robust accuracy varies with the amount of data used. For CIFAR-10 experiments in Figure 4c, the per-batch ratio is set to $0.5$, and $\beta$ is fixed at $0.4$. For SVHN in Figure 4d, the per-batch ratio is set to $0.5$, and $\beta$ is set to $0.6$. We aim to capture the robust accuracy at the optimal epoch in each case, and we observe that for Figures 4c and 4d, as the amount of data increases, the optimal accuracy is reached at a considerably later epoch.
(a) CIFAR-10
(b) SVHN
C-B Pre-Trained CLIP as Intermediate Model
In our experiments, we always began by training an intermediate model $f_{\hat{\theta}}$ using the same architecture as the final model $\theta_{\mathrm{final}}$. $f_{\hat{\theta}}$ was trained exclusively on the labeled dataset and then used to generate pseudo labels and select a subset from the extra dataset. However, when dealing with a generated or external dataset that already comes with labels, training an intermediate model specifically for selection undermines the data selection process’s intended benefits, primarily to reduce computational and time complexity. In application scenarios where the size of the labeled dataset $\mathcal{S}_{l}$ is also limited, likely, the intermediate model can not be trained with high accuracy, leading to large amounts of inaccurate pseudo labels.
To address these issues, we study the feasibility of using an alternative pre-trained model in LCS to perform data selection, specifically leveraging the CLIP model for extracting image embeddings [52]. CLIP, trained on a vast array of internet images and text, offers a rich and versatile feature space that can be utilized without additional training. By using the capabilities of pre-trained CLIP, we can bypass the intermediate step of training a model from scratch, thereby streamlining the selection process and further enhancing overall efficiency. This substitution allows us to select a relevant subset without having to train an intermediate model from scratch. Our experiments indicate that using CLIP for data selection yields results that are comparable to those obtained using the trained intermediate model, both in terms of standard and robust accuracies. Figure 8 shows our result when using CLIP instead of the previous intermediate model for data selection.
C-C SVHN with Ground-Truth Labels
We conduct additional experiments to test the performance of our data selection scheme on SVHN with pseudo labels replaced by ground-truth labels for the extra data. Table IV shows the comparison results. With ground-truth labels, our LCS schemes can achieve similar robust accuracy by selecting just $10\%$ extra data compared with no selection result.
C-D Resnet-18 Model Architecture
Our main experiments focus on the WideResNet architecture. We further evaluate the performance of SSAT under $\ell_{\infty}$ perturbations with $\epsilon=0.031$ using different data selection schemes and varying selection ratios with external unlabeled data on CIFAR-10 using a ResNet-18 model. Similar to our previous experiments, we train the ResNet-18 model on 5050K labeled samples from CIFAR-10 and a chosen subset of data from 500500K unlabeled images of Tiny-ImageNet. Our evaluation results are shown in Table V, which suggests that our proposed LCS methods can greatly improve the efficacy of SSAT on ResNet-18 while achieving comparable robustness performance of “No Selection” with $\alpha=100\%$.
TABLE V: Comparisons of SSAT performance on CIFAR-10 using the ResNet-18 architecture under various configurations. The model training dataset consists of 5050K labeled CIFAR-10 images and 500500K unlabeled images drawn from Tiny-ImageNet.
$\textbf{Ratio }\bm{\alpha}$ Method Clean ($\%$) PGD ($\%$)
@@ Appendix D Extended Analyses
TABLE VI: Performance of intermediate models trained using different schemes, including self-supervised learning (SSL), pre-trained CLIP, and fully-supervised methods, on CIFAR-10 under varying ratios of labeled data. We also present the corresponding performance of the final model produced by SSAT with LCS-KM, where the ratio of selected unlabeled data $\alpha$ is set at $10\%$.
@@
Training Method Clean Acc ($\%$) Clean Acc ($\%$) Robust Acc ($\%$)
@@ D-A Further Exploration of Generalizability
Perturbation Size. We consider external data from TinyImages as the unlabeled dataset. We investigate the effect of varying the $\ell_{\infty}$ perturbation size $\epsilon$, setting it to $4/255$ and $16/255$ while keeping TRADES as the adversarial training algorithm similar to the experiments in Table I. The results are depicted in Figures 9a and 9b, suggesting that our LCS-KM method is highly effective compared with random selection and no selection, almost approaching the robust accuracies achieved by full utilization of unlabeled data. The trends are shown to be consistent across varying perturbation sizes.
@@
Figure 10: Illustration of SSAT performance on CIFAR-10 under $\ell_{2}$ perturbations with $\epsilon=128/255$ using LCS-KM by varying $\alpha$: (a) external TinyImages and (b) DDPM-generated data.
$\bm{\ell}_{2}$ Perturbations. We further evaluate the performance of LCS-KM under $\ell_{2}$ perturbations with size $\epsilon=128/255$. In this experiment, we augment CIFAR-10 with either external unlabeled data sourced from Tiny-ImageNet or synthetic data generated via DDPM. We consider three configurations of selecting $10\%$, $20\%$, and $100\%$ extra unlabeled data. Our findings are shown in 10a and 10b, revealing that selecting $20\%$ unlabeled data achieves nearly equivalent robust accuracy compared to using the full dataset with $\alpha=100\%$. These results again confirm the efficiency and effectiveness of LCS-KM when adopted in SSAT algorithms.
@@ D-B Low Labeled Data Regime
We further study how the choice of the intermediate model and the amount of labeled data impact the final performance. For all experiments, we use a subset of the CIFAR-10 dataset as labeled data, supplemented with $10\%$ of data selected from a pool of 500500K unlabeled images sourced from Tiny-ImageNet. This additional data is selected using the intermediate model. The amount of unlabeled data remains constant across experiments, while the quantity of labeled data is varied. The intermediate model is trained using self-supervised learning (SSL) with a $5\%$ randomly selected subset of the unlabeled dataset, as long as the labeled data ratio is low. In addition, we include a pre-trained CLIP model [52] as an alternative intermediate model for comparison. The results of these experiments are presented in Table VI. The intermediate model’s quality significantly impacts the final model’s robust accuracy. In low-label regimes (e.g., $1\%$ or $10\%$), models trained using our intermediate model yield comparatively lower robust accuracy than those leveraging the pre-trained CLIP model. However, as the amount of labeled data for the intermediate model increases (e.g., 20% or 100%), the final model results are similar to or better than those of its counterparts based on the CLIP intermediate model. This trend demonstrates the effect of the intermediate model on the final robust accuracy. These observations underscore the importance of having a strong intermediate model for robust performance. Besides, we observe that pre-trained models like CLIP can be a good alternative to training an intermediate model, especially in low-labeled-data regimes. Detailed experiment results and discussions on the pre-trained CLIP intermediate model are provided in Appendix C-B.
D-C Further Discussion
Our work introduces several data reduction strategies that tackle the key challenges in adversarially robust learning, particularly the significant memory and computational overhead of state-of-the-art SSAT methods. By prioritizing boundary-adjacent data, these processes aim to enhance model robustness. However, these methods assume that the model’s latent representations are reliable for clustering and confidence evaluation, an assumption that may falter in the case of undertrained or poorly calibrated models. While these approaches effectively reduce the volume of unlabeled data required for training, their performance can be sensitive to hyperparameter configurations, such as the balance between boundary and non-boundary points and the size of the unlabeled dataset. By focusing on challenging examples near the decision boundary, these selection methods naturally address some aspects of robust overfitting and support robust generalization. Nevertheless, a deeper investigation is required to fully comprehend how these strategies influence decision boundary geometry and the model’s capacity for robust generalization. Combining these selection techniques with fast and efficient adversarial example generation methods, such as fast and free adversarial training, holds promise for further accelerating the training process while preserving robust performance.
Reducing the overall time required for adversarial training is critical for improving the efficiency and scalability of robust learning. Broadly, there are two complementary strategies to achieve this: (i) decreasing the computational cost of each training epoch, and (ii) reducing the total number of epochs needed for convergence. The first strategy aims to make each epoch faster without sacrificing robustness. Several prior works [18, 19, 53, 54] focus on this direction by introducing more efficient attack generation and optimization schemes. These approaches demonstrate that careful algorithmic design can substantially reduce per-epoch runtime, making adversarial training far more practical in large-scale settings. However, as shown in previous studies [15, 13], the total number of training epochs grows sharply with the size of the dataset. Consequently, even if the cost per epoch is reduced, increasing the amount of training data can still lead to a dramatic rise in overall training time. For example, adversarial training with approximately 11M generated samples typically requires around $400$ epochs to converge, whereas training with 100100M samples may demand as many as $2000$ epochs. Developing strategies that mitigate this dependence so that large-scale training does not necessitate a proportional increase in epochs would provide substantial computational savings and further improve the practicality of adversarially robust learning.

### v3

@@ C-B Pre-Trained CLIP as Intermediate Model
TABLE VI: Comparisons of SSAT performance on CIFAR-10 using the ResNet-18 architecture under various configurations. The model training dataset consists of 5050K labeled CIFAR-10 images and 500500K unlabeled images drawn from Tiny-ImageNet.
$\bm{\alpha}$ Method Clean ($\%$) PGD ($\%$)
@@
C-C SVHN with Ground-Truth Labels
We conduct additional experiments to test the performance of our data selection scheme on SVHN with pseudo labels replaced by ground-truth labels for the extra data. Table V shows the comparison results. With ground-truth labels, our LCS schemes can achieve similar robust accuracy by selecting just $10\%$ extra data compared with no selection result.
C-D ResNet-18 Model Architecture
Our main experiments focus on the WideResNet architecture. We further evaluate the performance of SSAT under $\ell_{\infty}$ perturbations with $\epsilon=0.031$ using different data selection schemes and varying selection ratios with external unlabeled data on CIFAR-10 using a ResNet-18 model. Similar to our previous experiments, we train the ResNet-18 model on 5050K labeled samples from CIFAR-10 and a chosen subset of data from 500500K unlabeled images of Tiny-ImageNet. Our evaluation results are shown in Table VI, which suggests that our proposed LCS methods can greatly improve the efficacy of SSAT on ResNet-18 while achieving comparable robustness performance of “No Selection” with $\alpha=100\%$.
Appendix D Extended Analyses
TABLE VII: Performance of intermediate models trained using different schemes, including self-supervised learning (SSL), pre-trained CLIP, and fully-supervised methods, on CIFAR-10 under varying ratios of labeled data. We also present the corresponding performance of the final model produced by SSAT with LCS-KM, where the ratio of selected unlabeled data $\alpha$ is set at $10\%$.
@@ Appendix D Extended Analyses
Training Method Clean Accuracy ($\%$) Clean Accuracy ($\%$) PGD Robust Accuracy ($\%$)
@@ D-A Further Exploration of Generalizability
Perturbation Size. We consider external data from TinyImages as the unlabeled dataset. We investigate the effect of varying the $\ell_{\infty}$ perturbation size $\epsilon$, setting it to $4/255$ and $16/255$ while keeping TRADES as the adversarial training algorithm, similar to the experiments in Table I. The results are depicted in Figures 8a and 8b, suggesting that our LCS-KM method is highly effective compared to random selection and no selection, approaching the robust accuracies achieved by fully utilizing unlabeled data. The trends are consistent across varying perturbation sizes.
@@
Figure 9: Illustration of SSAT performance on CIFAR-10 under $\ell_{2}$ perturbations with $\epsilon=128/255$ using LCS-KM by varying $\alpha$ with: (a) 500500K external TinyImages and (b) 11M DDPM-generated data.
$\bm{\ell}_{2}$ Perturbations. We further evaluate the performance of LCS-KM under $\ell_{2}$ perturbations with size $\epsilon=128/255$. In this experiment, we augment CIFAR-10 with either external unlabeled data sourced from Tiny-ImageNet or synthetic data generated via DDPM. We consider three configurations of selecting $10\%$, $20\%$, and $100\%$ extra unlabeled data. Our findings are presented in Figures 9a and 9b, which reveal that selecting $20\%$ unlabeled data achieves nearly equivalent robust accuracy compared to using the full dataset with $\alpha=100\%$. These results again confirm the efficiency and effectiveness of LCS-KM when adopted in SSAT algorithms.
@@ D-B Low Labeled Data Regime
We further investigate how the choice of intermediate model and the amount of labeled data affect the final performance. For all experiments, we use a subset of the CIFAR-10 dataset as labeled data, supplemented with $10\%$ of data selected from a pool of 500500K unlabeled images sourced from Tiny-ImageNet. This additional data is selected using the intermediate model. The amount of unlabeled data remains constant across experiments, while the quantity of labeled data is varied. The intermediate model is trained using self-supervised learning (SSL) with a $5\%$ randomly selected subset of the unlabeled dataset, provided the labeled data ratio is low. In addition, we include a pre-trained CLIP model [32] as an alternative intermediate model for comparison.
Table VII summarizes our results. The quality of the intermediate model significantly impacts the robust accuracy of the final model. In low-label regimes (e.g., $1\%$ or $10\%$), models trained using our intermediate model yield comparatively lower robust accuracy than those leveraging the pre-trained CLIP model. However, as the amount of labeled data for the intermediate model increases (e.g., $20\%$ or $100\%$), the performance of the final model becomes similar to or better than that of its counterparts based on the CLIP intermediate model. This trend demonstrates the effect of the intermediate model on the final robust accuracy. These observations highlight the importance of having a robust intermediate model for achieving strong performance. Additionally, we observe that pre-trained models, such as CLIP, can be a good alternative to training an intermediate model, especially in low-labeled-data regimes. Detailed results and discussions on the pre-trained CLIP intermediate model are provided in Appendix C-B.
TABLE VIII: Full version of Table IV. We evaluate full SSAT ($\alpha=100\%$) on CIFAR-10 under $\ell_{\infty}$ perturbations with $\epsilon=0.031$ under various learning rate schedulers with/without early stopping.
Setting Schedule Epoch Clean ($\%)$ PGD ($\%)$ AA ($\%)$
External Cosine $100$ $84.1$ $54.0$ $50.4$
$200$ $89.3$ $62.2$ $59.8$
Stepwise $100$ $88.0$ $60.6$ $57.6$
$200$ $89.7$ $62.5$ $58.6$
DDPM Cosine $100$ $78.6$ $52.1$ $48.6$
$400$ $86.2$ $62.4$ $59.9$
Stepwise $100$ $83.0$ $58.5$ $55.9$
$400$ $85.7$ $61.8$ $58.4$
D-C Performance of Full SSAT Stopped at Earlier Epochs
To investigate whether early stopping can serve as an effective strategy for full SSAT rather than training for the complete number of epochs, we evaluate SSAT models on CIFAR-10 using $100\%$ of the unlabeled data with two popular learning rate scheduling schemes, both with an initial learning rate of $0.1$: (Cosine) cosine annealing and (Stepwise) stepwise decay with a decaying factor of $10$ at training epochs $75$ and $90$. We compare the performance of SSAT stopped at epoch $100$, the same as our data reduction methods, against SSAT models trained for the full duration: $200$ epochs for 500500K external TinyImages and $400$ epochs for 11M DDPM-generated data. Table VIII presents the complete results. Below, we summarize the key observations regarding the interplay between learning rate scheduling and the training duration of SSAT:
Cosine vs. Stepwise. The two learning rate schedulers exhibit distinct training dynamics during the early epochs. At epoch $100$, models trained with cosine annealing show significantly lower performance (e.g., $50.4\%$ AA robust accuracy for external data and $48.6\%$ for DDPM-generated data) compared to stepwise decay ($57.6\%$ and $55.9\%$, respectively). However, when trained to completion, cosine annealing achieves more competitive final robustness performance than stepwise scheduling, approximately $1\%-1.5\%$ improvement in robust accuracy when evaluated using AutoAttack.
Effectiveness of Early Stopping. For the stepwise scheduler, early stopping at epoch $100$ can achieve robust accuracy not far from the robustness of the final model returned by full SSAT. For external data, the robustness gap between epochs $100$ and $200$ is only $1\%$ with a doubled training time. Similarly, for DDPM data, early stopping at epoch $100$ yields $55.9\%$ AA robust accuracy compared to $58.4\%$ at epoch $400$, a difference of $2.5\%$ with $4\times$ more training. In comparison, early stopping is not effective for full SSAT when using the cosine annealing scheduler. Prolonged training of full SSAT algorithms from epoch $100$ gradually increases the clean and robust accuracy performance until reaching peak performance, regardless of the type of learning rate scheduler implemented. These observations are aligned with the full SSAT curves illustrated in Figure 1d, suggesting that learning is not complete for SSAT with $\alpha=100\%$ if it is simply stopped at an early epoch.
Comparison with LCS-KM. Even with the stepwise scheduler and early stopping employed, full SSAT early stopped at epoch $100$ achieves AA robust accuracy of $57.6\%$ (external) and $55.9\%$ (DDPM-generated), slightly lower than our LCS-KM method with $20\%$ unlabeled data that achieves $57.8\%$ and $57.2\%$ for the respective settings (Table I). Under cosine scheduling, the robustness improvement of using LCS-KM is much more pronounced compared to full SSAT with early stopping. These results demonstrate the advantages of our data reduction methods, offering a viable solution to improve the data efficiency of SSAT algorithms and accelerate the learning process while largely preserving their robustness advantages.
D-D Further Discussions
Our work introduced several strategies to improve the data efficiency of SSAT algorithms by prioritizing unlabeled data adjacent to the model’s decision boundaries. However, the success of our methods hinges on the hypothesis that the intermediate model’s latent representations are reliable for pseudo-labeling and prioritizing unlabeled data, thereby improving robust generalization, which may falter in scenarios where the model is undertrained or poorly calibrated. In addition, while our methods effectively reduce the volume of unlabeled data, their final performance can be sensitive to hyperparameter configurations, such as the balance between boundary and non-boundary points. Developing automatic tools to automate the process of hyperparameter selection, as well as conducting in-depth theoretical studies to fully comprehend how these strategies influence decision boundary geometry and the model’s capacity for robust generalization, along with their adaptability to potential variations, would be meaningful future directions.
Moreover, reducing the overall runtime required for adversarial training without sacrificing robustness is critical for achieving efficient and scalable robust learning. In the prior literature, two complementary strategies have been proposed: (i) decreasing the computation of each training epoch, such as free and fast adversarial training [37, 45], and other more efficient optimizers for generating perturbations in an adversarial training framework [7, 53], and (ii) reducing the total number of epochs needed for convergence, such as implementing early stopping [33]. While the former techniques can reduce the per-epoch runtime, the overall training process is less stable compared to vanilla adversarial training, where PGD with sufficient iterations is used to solve the inner maximization problem. Our work falls into the second category, where we adopt early stopping to accelerate the convergence of SSAT algorithms. Nevertheless, developing effective strategies based on our data reduction schemes and more efficient inner maximization solvers represents a promising future direction for further improving the practicality of SSAT algorithms.